OpenStudy (anonymous):

3/x+5 - 3x/3x+15 + 1/2x+10

OpenStudy (mathteacher1729):

Two questions: 1) Is this the expression? \[\frac{3}{x+5}-\frac{3x}{3+15}+\frac{1}{2x+10}\]2) Wht do they want you to do here? (I'm guessing they want you to simplify it into a single fraction...)

OpenStudy (anonymous):

find the lcd

OpenStudy (anonymous):

then simplify the answer

OpenStudy (anonymous):

look 3x/3x= zero so just get rid of that then 5 -15= -10 -10 +10 = 0 so ur left with 3/x + 1/2x which is just 4/x

OpenStudy (angela210793):

3x/3x doesn't equal 0 but 1....O.o

OpenStudy (anonymous):

oh yh whopps so its just 4/x +x

OpenStudy (anonymous):

or simplified 6/x

OpenStudy (angela210793):

Don't u have to find the common denom. in here?

OpenStudy (anonymous):

wrong the answer is 7-2x/2(x+5)

OpenStudy (angela210793):

U found it by urself....gr8 :)

OpenStudy (anonymous):

you have to find the lcd

OpenStudy (angela210793):

wht is lcd?